# Privacy-Preserving Instructions for Aligning Large Language Models

### v1

@@ 4.2 Measuring the Distributional Gap
We compute the MAUVE score (Pillutla et al., 2022) between synthetic and real instructions. MAUVE compares two text distributions using divergence frontiers (Kynkäänniemi et al., 2019; Djolonga et al., 2020) and has been widely used to analyze the gap between synthetic text and real text (Yue et al., 2022; Kurakin et al., 2023). We use two representations of text distributions to compute the MAUVE scores: the unigram distribution, and the distribution of quantized text embeddings (Pillutla et al., 2021). Throughout the paper, we compute embeddings of instructions using the Sentence-T5-base model (Ni et al., 2021). Though the absolute values of MAUVE depend on the setup used to calculate it, such some scalar scaling constant, the relative ranking of different datasets is stable. Our setup for computing MAUVE scores is in Appendix F.1.
Table 1: MAUVE scores with Chatbot Arena instructions as the target dataset. Best scores of DP models are marked in bold. Private synthetic instructions achieves significantly better scores than non-private but out-of-distribution FLAN instructions. Furthermore, filtering synthetic instructions with Algorithm 2 improves the scores by large margins.
@@
Figure 5: Running Algorithm 2 with different $K$ and $\sigma$. The MAUVE scores initially improve with an increase in $K$, then they either start to plateau or decline. Concurrently, the number of required initial samples for selecting a subset of 180 thousand increases with $K$.
Table 1 presents the main MAUVE scores. We compute the scores with three random seeds and report the average. The first observation is that the MAUVE scores of FLAN instructions are notably lower than the scores of (private) synthetic instructions, indicating that there is a significant distributional gap between FLAN instructions and user queries. Additionally, filtering synthetic instructions with Algorithm 2 improves MAUVE scores by large margins.
We also run Algorithm 2 with different values of $K$ and $\sigma$. The resulting MAUVE scores and the required quantity of initial synthetic instructions are presented in Figure 5. The number of clustering bins considered ranges from 100 to 2000. The value of $\sigma$ is chosen from $[10,30]$. As shown in Figure 5, the MAUVE scores initially improve with an increase in $K$. However, the scores begin to either plateau or decline for large values of $K$. This is likely because a larger value of $K$ weakens the aggregated counts, rendering the filtering process more sensitive to noise. Interestingly, the number of initial synthetic instructions required by Algorithm 2 to select a subset of 180 thousand also increases with $K$. In the rest of this paper, we set $K=1000$ and $\sigma=10$ for filtering one million initial synthetic instructions. In Appendix D.3, we run Algorithm 2 with fewer initial synthetic instructions. The results suggest that filtering 500 thousand initial synthetic instructions results in inferior MAUVE scores compared to using one million initial instructions for large values of $K$. Notably, to select a subset of 180 thousand from 500 thousand initial samples, we have to use sampling with replacement in the re-sampling process for large values of $K$.
@@ 4.3 Supervised Fine-tuning
Table 2: Win-rates evaluated on Chatbot Arena instructions (over 3 seeds). The left column shows the fine-tuning sets (default size is 180K), except for Vicuna-v1.3 which is already instruction-tuned and is not further fine-tuned. The models trained with filtered private synthetic instructions are comparable to those trained with real instructions, and improve over baselines by large margins.
7B Models FLAN (non-private) $50\%$ Vicuna-v1.3 $64.1\%$ ($\pm 0.61$) Chatbot Arena (non-private) $68.9\%$ ($\pm 0.31$) Chatbot Arena ($\varepsilon=5.94$) $60.7\%$ ($\pm 0.42$) Synthetic ($\varepsilon=5.94$, no filt.) $62.7\%$ ($\pm 0.34$) Synthetic ($\varepsilon=5.98$) $67.8\%$ ($\pm 0.32$) Synthetic (300K, $\varepsilon=5.98$) 68.1% ($\pm 0.37$) 13B Models Vicuna-v1.3 $72.8\%$ ($\pm 0.58$) Synthetic (300K, $\varepsilon=5.98$) 74.5% ($\pm 0.41$)
@@
Table 2 presents the win-rates of evaluating various models on 800 random instructions from the Chatbot Arena test split. The baseline model is LLaMA 7B fine-tuned with 180 thousand FLAN instructions. We also include Vicuna-v1.3 (Chiang et al., 2023) in our evaluation, one of the state-of-the-art models that is non-privately fine-tuned from LLaMA on conversations collected from the ShartGPT website (ShareGPT, 2023). The results in Table 6 suggest two main findings. First, private synthetic instructions generated by our algorithms are of high utility, only showing a minor performance drop when compared to using real instructions without any privacy protection. Second, fine-tuning the models on the domain of private instructions yields significantly better performance compared to fine-tuning the models on out-of-domain FLAN instructions. In addition to Chatbot Arena instructions, we also evaluate the models on the AlpacaEval benchmark (Dubois et al., 2023), which comprises 805 instructions collected from public evaluation sets. We present the results in Appendix D.4. It is noteworthy that AlpacaEval instructions do not accurately represent real-world user queries. Nonetheless, such evaluation allows us to make direct comparisons with models listed on the AlpacaEval leaderboard (Li et al., 2023a). When evaluated on AlpacaEval, the best performing DP models slightly underperform Vicuna-v1.3 but are still comparable with other leading opensource models such as LLaMA2-Chat (Touvron et al., 2023b).
4.4 RLHF with Proximal Policy Optimization
@@ 5 Conclusion
For the common practice of fine-tuning LLMs on instructions, we identify a new threat model on the user provided instructions: potential risks in exposing sensitive personal information in the instructions to human annotators. To address this concern, as well as preventing fine-tuned LLMs to memorize sensitive information in user instructions, we propose to replace user instructions with DP synthetic instructions and introduce a novel framework for generating high-quality instructions by using a DP histogram to resample synthetic instructions to match the distribution of the real dataset in the embedding space. The effectiveness of our framework is validated through comprehensive experiments on publicly available Chatbot Arena datasets (Zheng et al., 2023a; b) and LLaMA and Phi-1.5 models (Touvron et al., 2023a; Li et al., 2023b). Important future directions include developing algorithms capable of privately learning from user instructions in multiple modalities, including image and speech.
@@ Appendix A Additional Related Work
Wu et al. (2023) present pioneering efforts in aligning language models with differential privacy. They align GPT-2 models (Radford et al., 2019) using private optimizers to ensure that models trained with both supervised fine-tuning and reinforcement learning satisfy differential privacy. They demonstrate that the rewards achieved by DP models are comparable to those of non-private models on the IMDb dataset (Maas et al., 2011) and the Reddit TL;DR dataset (Völske et al., 2017). Compared to Wu et al. (2023), this work introduces several new aspects. First, our framework mitigates privacy risks arising from both human annotation and model memorization. In contrast, the privacy guarantees in Wu et al. (2023) only apply to the trained models. Second, our experiments focus on instruction datasets collected during the real-world deployment of LLMs (Zheng et al., 2023a). Moreover, we evaluate the aligned models’ ability to follow instructions not only through rewards but also through win-rates evaluated by LLM judges (Zheng et al., 2023c; Dubois et al., 2023).
@@
Lin et al. (2023) investigate generating realistic differentially private synthetic images using only API access to foundation models. Their approach, termed Private Evolution, also leverages votes from real samples to select synthetic samples. Private Evolution is a multi-round generation algorithm. In each round, it first generates a pool of candidate images by using variation APIs to create variations for samples from the previous round. Then they sample (with replacement) from the candidate pool based on votes from real images. The key difference between their selection process with ours is that we cluster synthetic samples to improve the signal-to-noise ratio during the selection process. Our Algorithm 2 allows us to run the selection process with a minimal privacy cost. Additionally, we focus on training instruction-following models and conduct comprehensive experiments to verify the effectiveness of our algorithms.
@@
Recently, Zhang et al. (2023b) show that users may disclose private information in their conversations with LLMs. They investigate the conversations collected on the ShareGPT website. In this work, we investigate the user instructions in the Chatbot Arena Conversations dataset [^1] https://huggingface.co/datasets/lmsys/chatbot_arena_conversations. Compared to the analysis in Zhang et al. (2023b), our analysis is new in two aspects. First, all conversations from ShareGPT are proactively shared by the user themselves and hence may result in an inherent sample bias. In contrast, users instructions from Chatbot Arena are directly collected by the service provider. Second, we conduct a quantitative analysis of six distinct categories of personally identifiable information, providing a more comprehensive understanding for the presence of common types of sensitive information.
@@
In addition to DP synthetic text, significant research has been dedicated to generating other forms of data with differential privacy, including synthetic tabular data (Tantipongpipat et al., 2019; Ge et al., 2020; Tao et al., 2021; Rosenblatt et al., 2020; Aydore et al., 2021; Vietri et al., 2020; 2022; McKenna et al., 2021; 2022; Liu et al., 2023) and synthetic image data (Xie et al., 2018; Torkzadehmahani et al., 2019; Harder et al., 2021; Dockhorn et al., 2022; Ghalebikesabi et al., 2023; Sehwag et al., 2023). In the realm of generating DP synthetic tabular data, recent studies have explored enhancing the quality of synthetic data through post-processing the initial outputs (Neunhoeffer et al., 2020; Liu et al., 2021; Wang et al., 2023b). Most recently, Wang et al. (2023b) propose to privatize the correlation matrix of a private dataset, and then re-sample the initial synthetic data based on the privatized correlation matrix. Their approach achieves state-of-the-art performance on several benchmarks. To the best of our knowledge, this work marks the first attempt to investigate the post-processing of private synthetic text. We introduce a filtering algorithm that demonstrates both high effectiveness and minimal additional privacy costs.
@@
Leveraging pre-trained models in private learning has substantially improved the privacy-utility trade-off of deep learning with DP. This advancement extends to various domains, including natural language processing (Majmudar et al., 2022; Bu et al., 2023; Wang et al., 2023a; Du et al., 2023; Xiao et al., 2023; Duan et al., 2023; Tang et al., 2023b; 2024; Panda et al., 2023; Hong et al., 2023), computer vision (Luo et al., 2021; Golatkar et al., 2022; Mehta et al., 2022; Panda et al., 2022; De et al., 2022; Zhu et al., 2023; Tang et al., 2023a; Xu et al., 2023; Thaker et al., 2023; Zhang et al., 2023a), and speech recognition (Shamsabadi et al., 2022; Azam et al., 2023). Recent theoretical studies have provided evidence to explain the efficacy of using pre-trained models in private learning (Li et al., 2022a; Ganesh et al., 2023). In this paper, we follow this line of research and utilize DP fine-tuning to train generative models for creating synthetic instructions. It is noteworthy that if the pre-training process does not satisfy differential privacy, then DP fine-tuning only assures the privacy of fine-tuning data. In real-world deployments, it is important to also consider the privacy risks of pre-training data (Tramèr et al., 2022). One potential solution is to also enforce differential privacy in the pre-training stage (Kurakin et al., 2022; Anil et al., 2022; Sander et al., 2022).
@@ Appendix B Real-world User Instructions Are Sensitive
Figure 8: The prompt for initial screening with GPT-3.5-Turbo. The categories of sensitive texts are inspired by the categories of personally identifiable information in Pilán et al. (2022). The authors label two emails from a public dataset as demonstrations (see Figure 16 for one of them).
@@ Appendix C Empirical Privacy Leakage
Figure 12: Six canary instructions. Each canary contains either a fake phone number or a fake address. Canaries are repeated with 1, 10, or 100 times. Phone numbers are randomized 10-digit sequences, and addresses are generated using an off-the-shelf tool.
@@
One common empirical approach to assess the privacy risk of a training algorithm is injecting canaries (specifically crafted training samples) into the training set, and evaluating the extent to which the canaries are exposed through querying the trained model (Pillutla et al., 2023; Nasr et al., 2023b; Jagielski, 2023). In this section, we follow this approach to compare the empirical privacy leakage of models trained with/without differential privacy. We use six canary instructions, three of them contain phone numbers and the other three contain addresses (see Figure 12 for the six canaries). Following Yue et al. (2022), we repeat the canaries with different times (1, 10, and 100) to explore the worst case privacy leakage (Lee et al., 2021; Kandpal et al., 2022).
Table 3: Measuring privacy leakage using canaries. ’Loss Rank’ is the rank of loss among 10 thousand random canaries, taking the smaller rank of the two types of canaries. The rightmost column shows how many times the phone numbers and addresses in the injected canaries appear in one million synthetic instructions.
@@
Repetition $\varepsilon$ Loss Rank # of Occurrences 1 non-private 50 0, 0 $\varepsilon=2.86$ 4142 0, 0 $\varepsilon=5.94$ 3566 0, 0 10 non-private 1 2, 0 $\varepsilon=2.86$ 1232 0, 0 $\varepsilon=5.94$ 1254 0, 0 100 non-private 1 605, 290 $\varepsilon=2.86$ 806 0, 0 $\varepsilon=5.94$ 698 0, 0
Table 3 presents the empirical leakage of canary instructions. The exposure of these canaries is quantified using two simple measures. The first one is the relative position of the loss of a inserted canary among the losses of random canaries. We generate 10 thousand random canaries by replacing the phone numbers and addresses with random ones while keep the other parts unchanged. The second one is the frequency of occurrences of the inserted phone numbers and addresses within the synthetic instructions generated by a trained model. We generate one million synthetic instructions for each trained model, which is more than five times larger than the training set. The results in Table 3 suggest that for generative models trained without differential privacy, the loss ranks of all inserted canaries are below 100. Furthermore, in scenarios where canaries are replicated 10 or 100 times, the synthetic instructions leak the inserted phone numbers or/and addresses. In contrast, for models trained with differential privacy, the loss ranks of all inserted canaries are significantly higher, and no leakage of the inserted numbers and addresses is observed.

### v2

@@ 4.2 Measuring the Distributional Gap
We compute the MAUVE score (Pillutla et al., 2021, 2022) between synthetic and real instructions. MAUVE compares two text distributions using divergence frontiers (Kynkäänniemi et al., 2019; Djolonga et al., 2020) and has been widely used to analyze the gap between synthetic text and real text (Yue et al., 2022; Kurakin et al., 2023). We use two representations of text distributions to compute the MAUVE scores: the unigram distribution, and the histogram of clustered text embeddings (Pillutla et al., 2021). Our setup for computing MAUVE scores is in Appendix F.1.
Table 1: MAUVE scores with Chatbot Arena instructions as the target dataset. Best scores of DP models are marked in bold. Filtering synthetic instructions with Algorithm 2 improves the scores of private synthetic instructions by large margins.
@@
Figure 5: Running Algorithm 2 with different $K$ and $\sigma$. The MAUVE scores initially improve with an increase in $K$, then they either start to plateau or decline.
Table 1 presents the main MAUVE scores. We compute the scores with three random seeds and report the average. Filtering private synthetic instructions with Algorithm 2 improves MAUVE scores by large margins. We further demonstrate improvements in MAUVE scores for two other datasets, IMDB movie reviews and PubMed abstracts, to underscore the wide applicability of Algorithm 2. Details on the results of the two additional datasets are in Appendix D.4. Additionally, there is a significant distributional gap between FLAN instructions and real-world user instructions. This indicates that in-domain user instructions, while being sensitive, play a crucial role in LLM alignment and should be utilized in a privacy-preserving way.
Figure 6: The number of required initial samples for selecting a subset of 180 thousand increases with $K$.
We also run Algorithm 2 with different values of $K$ and $\sigma$. The values of $K$ ranges from 100 to 2000, and the values of $\sigma$ are chosen from $\{10,30\}$. The resulting MAUVE scores are presented in Figure 5. The required quantity of initial synthetic instructions as $K$ varies is presented in Figure 6. The MAUVE scores initially improve with an increase in $K$. However, the scores begin to either plateau or decline for large values of $K$. This is likely because a larger value of $K$ weakens the aggregated counts, rendering the filtering process more sensitive to noise. Interestingly, the number of initial synthetic instructions required by Algorithm 2 to select a subset of 180 thousand also increases with $K$. In the rest of this paper, we set $K=1000$ and $\sigma=10$ for filtering one million initial synthetic instructions. In Appendix D.3, we run Algorithm 2 with fewer initial synthetic instructions. The results suggest that filtering 500 thousand initial synthetic instructions results in inferior MAUVE scores compared to using one million initial instructions. Notably, to select a subset of 180 thousand from 500 thousand initial samples, we have to use sampling with replacement for large values of $K$.
@@ 4.3 Supervised Fine-tuning
Table 2: Win-rates on Chatbot Arena instructions over 3 seeds. The left column shows the fine-tuning sets (default size is 180K), except for Vicuna-v1.3 which is a public model and is not further fine-tuned. Best win-rates of DP models are in bold. The win rate with filtering ($\varepsilon=2.91$) surpasses that without filtering ($\varepsilon=5.94$).
7B Models FLAN (non-private) $50\%$ Vicuna-v1.3 $64.1\%$ ($\pm 0.61$) Chatbot Arena (non-private) $68.9\%$ ($\pm 0.31$) Chatbot Arena ($\varepsilon=5.94$) $60.7\%$ ($\pm 0.42$) Synthetic ($\varepsilon=5.94$, no filt.) $62.7\%$ ($\pm 0.34$) Synthetic ($\varepsilon=2.91$) $65.6\%$ ($\pm$0.36) Synthetic ($\varepsilon=5.98$) $67.8\%$ ($\pm 0.32$) Synthetic (300K, $\varepsilon=5.98$) 68.1% ($\pm 0.37$) 13B Models Vicuna-v1.3 $72.8\%$ ($\pm 0.58$) Synthetic ($\varepsilon=5.94$, no filt.) $71.5\%$ ($\pm$0.40) Synthetic ($\varepsilon=2.91$) $73.2\%$ ($\pm$0.48) Synthetic ($\varepsilon=5.98$) $74.0\%$ ($\pm 0.62$) Synthetic (300K, $\varepsilon=5.98$) 74.5% ($\pm 0.41$)
Table 2 presents the win-rates for various models evaluated on 800 random instructions from the Chatbot Arena test split. The baseline model is LLaMA 7B fine-tuned with 180 thousand FLAN instructions. We also run evaluations using the baseline model trained on non-private Chatbot Arena data, with the results presented in Appendix D.5. The findings are consistent with those in Table 2. We include Vicuna-v1.3 (Chiang et al., 2023) in our evaluation, one of the state-of-the-art models that is non-privately fine-tuned from LLaMA on conversations collected from the ShartGPT website (ShareGPT, 2023). The results in Table 2 suggest two main findings. First, private synthetic instructions generated by our algorithms are of high utility, only showing a minor performance drop when compared to using real instructions without any privacy protection. Second, fine-tuning the models on the domain of private instructions yields significantly better performance compared to fine-tuning the models on out-of-domain FLAN instructions. In addition to Chatbot Arena instructions, we also evaluate the models on the AlpacaEval benchmark (Dubois et al., 2023), which comprises 805 instructions collected from public evaluation sets. We present the results in Appendix D.6. It is noteworthy that AlpacaEval instructions do not accurately represent real-world user queries. Nonetheless, such evaluation allows us to make direct comparisons with models listed on the AlpacaEval leaderboard (Li et al., 2023a). When evaluated on AlpacaEval, the best performing DP models slightly underperform Vicuna-v1.3 but are still comparable with other leading opensource models such as LLaMA2-Chat (Touvron et al., 2023b).
4.4 RLHF with Proximal Policy Optimization
@@ 5 Conclusion
For the common practice of fine-tuning LLMs on instructions, we identify a new threat model on the user provided instructions: potential risks in exposing sensitive personal information in the instructions to human annotators. To address this concern, as well as preventing fine-tuned LLMs to memorize sensitive information in user instructions, we propose to replace user instructions with DP synthetic instructions and introduce a novel framework for generating high-quality instructions by using a DP histogram to resample synthetic instructions to match the distribution of the real dataset in the embedding space. The effectiveness of our framework is validated through comprehensive experiments on publicly available Chatbot Arena datasets (Zheng et al., 2023a, b) and LLaMA and Phi-1.5 models (Touvron et al., 2023a; Li et al., 2023b). Important future directions include developing algorithms capable of privately learning from user instructions in multiple modalities, including image and speech.
@@ Acknowledgement
Impact Statement
In Appendix B, we describe a method for identifying sensitive information within instructions. Our goal is to demonstrate the vulnerability of real-world user inputs, in the hope of encouraging leading AI companies to prioritize the protection of user privacy. While it is possible that potential adversaries might employ the described method, we argue that the negative societal impact of this work is minimal. The method we employ is standard, and the instructions under analysis are already public.
@@ Appendix A Additional Related Work
Wu et al. (2024a) present pioneering efforts in aligning language models with differential privacy. They align GPT-2 models (Radford et al., 2019) using private optimizers to ensure that models trained with both supervised fine-tuning and reinforcement learning satisfy differential privacy. They demonstrate that the rewards achieved by DP models are comparable to those of non-private models on the IMDb dataset (Maas et al., 2011) and the Reddit TL;DR dataset (Völske et al., 2017). Compared to Wu et al. (2024a), this work introduces several new aspects. First, our framework mitigates privacy risks arising from both human annotation and model memorization. In contrast, the privacy guarantees in Wu et al. (2024a) only apply to the trained models. Second, our experiments focus on instruction datasets collected during the real-world deployment of LLMs (Zheng et al., 2023a). Moreover, we evaluate the aligned models’ ability to follow instructions not only through rewards but also through win-rates evaluated by LLM judges (Zheng et al., 2023c; Dubois et al., 2023).
@@
Lin et al. (2023) and Xie et al. (2024) investigate generating differentially private synthetic data using only API access to foundation models. Their approach, termed Private Evolution, also leverages votes from real samples to select synthetic samples. Private Evolution is a multi-round generation algorithm. In each round, it first generates a pool of candidate synthetic samples by applying variation APIs to samples from the previous round. Then they sample from the candidate pool based on votes from real datapoints. The key difference between their selection process with ours is that we cluster synthetic samples to improve the signal-to-noise ratio during the selection process. Our Algorithm 2 allows us to run the selection process with a minimal privacy cost. Additionally, we focus on training instruction-following models and conduct comprehensive experiments to verify the effectiveness of our algorithms.
@@
Recently, Zhang et al. (2023b) show that users may disclose private information in their conversations with LLMs. They investigate the conversations collected on the ShareGPT website. In this work, we investigate the user instructions in the Chatbot Arena Conversations dataset [^2] https://huggingface.co/datasets/lmsys/chatbot_arena_conversations.. Compared to the analysis in Zhang et al. (2023b), our analysis is new in two aspects. First, all conversations from ShareGPT are proactively shared by the user themselves and hence may result in an inherent sample bias. In contrast, users instructions from Chatbot Arena are directly collected by the service provider. Second, we conduct a quantitative analysis of six distinct categories of personally identifiable information, providing a more comprehensive understanding for the presence of common types of sensitive information.
@@
In addition to DP synthetic text, significant research has been dedicated to generating other forms of data with differential privacy, including synthetic tabular data (Jordon et al., 2018; Tantipongpipat et al., 2019; Ge et al., 2020; Tao et al., 2021; Rosenblatt et al., 2020; Aydore et al., 2021; Vietri et al., 2020, 2022; McKenna et al., 2021, 2022; Liu et al., 2023) and synthetic image data (Xie et al., 2018; Torkzadehmahani et al., 2019; Harder et al., 2021; Dockhorn et al., 2022; Ghalebikesabi et al., 2023; Sehwag et al., 2023). In the realm of generating DP synthetic tabular data, recent studies have explored enhancing the quality of synthetic data through post-processing the initial outputs (Neunhoeffer et al., 2020; Liu et al., 2021; Wang et al., 2023b). Most recently, Wang et al. (2023b) propose to privatize the correlation matrix of a private dataset, and then re-sample the initial synthetic data based on the privatized correlation matrix. Their approach achieves state-of-the-art performance on several benchmarks. To the best of our knowledge, this work marks the first attempt to investigate the post-processing of private synthetic text. We introduce a filtering algorithm that demonstrates both high effectiveness and minimal additional privacy costs.
@@
Leveraging pre-trained models in private learning has substantially improved the privacy-utility trade-off of deep learning with DP. This advancement extends to various domains, including natural language processing (Majmudar et al., 2022; Bu et al., 2023; Wang et al., 2023a; Du et al., 2023; Xiao et al., 2023; Duan et al., 2023a; Tang et al., 2023b, 2024a; Panda et al., 2023; Hong et al., 2023), computer vision (Luo et al., 2021; Golatkar et al., 2022; Mehta et al., 2022; Panda et al., 2022; De et al., 2022; Zhu et al., 2023; Tang et al., 2023a; Xu et al., 2023; Thaker et al., 2023; Zhang et al., 2023a; Tobaben et al., 2023), and speech recognition (Shamsabadi et al., 2022; Azam et al., 2023). Recent theoretical studies have provided evidence to explain the efficacy of using pre-trained models in private learning (Li et al., 2022a; Ganesh et al., 2023). In this paper, we follow this line of research and utilize DP fine-tuning to train generative models for creating synthetic instructions. It is noteworthy that if the pre-training process does not satisfy differential privacy, then DP fine-tuning only assures the privacy of fine-tuning data. In real-world deployments, it is important to also consider the privacy risks of pre-training data (Tramèr et al., 2022). One potential solution is to also enforce differential privacy in the pre-training stage (Kurakin et al., 2022; Anil et al., 2022; Sander et al., 2022).
@@ Appendix B Real-world User Instructions Are Sensitive
Figure 9: The prompt for initial screening with GPT-3.5-Turbo. The categories of sensitive texts are inspired by the categories of personally identifiable information in Pilán et al. (2022). The authors label two emails from the Enron Email dataset as demonstrations.
@@ Appendix C Empirical Privacy Leakage
Figure 13: Six canary instructions. Each canary contains either a fake phone number or a fake address.
@@
One common empirical approach to assess the privacy risk of a training algorithm is injecting canaries (specifically crafted training samples) into the training set, and evaluating the extent to which the canaries are exposed through querying the trained model (Pillutla et al., 2023; Nasr et al., 2023b; Jagielski, 2023). In this section, we follow this approach to compare the empirical privacy leakage of models trained with and without DP. We use six canary instructions, three of them contain phone numbers and the other three contain addresses (see Figure 13 for the six canaries). Phone numbers are randomized 10-digit sequences, and addresses are generated using an off-the-shelf tool [^3] https://pypi.org/project/random-address.. Following Yue et al. (2022), we repeat the canaries with different times (1, 10, and 100) to explore the worst case privacy leakage (Lee et al., 2021; Kandpal et al., 2022). It is important to note that the privacy parameter $\varepsilon$ for a canary linearly increases with the number of repetitions (Dwork et al., 2014; Vadhan, 2017).
Table 3: Evaluating the leakage of injected canaries. ‘Loss Rank’ is the rank of a canary among 10,000 instructions that use the same template but different secrets. The rightmost two columns indicate if the trained model leaks the secrets during inference, conducted in two ways. First, we generate 100,000 synthetic instructions without prompts and check for the presence of secrets. Second, we use text preceding the secret as prompt to see if the model’s greedy decoding leaks the secret.
@@
Instance-level $\varepsilon$ Repetition Loss Rank Leakage (No Prompt) Leakage (With Prompt) Non-private 1 50 0/2 0/2 10 1 1/2 2/2 100 1 2/2 2/2 $\varepsilon=2.86$ 1 4142 0/2 0/2 10 1232 100 806 $\varepsilon=5.94$ 1 3566 0/2 0/2 10 1254 100 698
We quantify the leakage using two measures. The first measure is the relative position of a canary’s loss among the losses of 10,000 instructions. These instructions follow the same template but differ in the phone numbers or addresses they contain. For each repetition count, we calculate two loss ranks: one for a canary with a phone number and another for a canary with an address. We only report the lower of these two ranks. The second measure evaluates whether the trained model leaks the secret verbatim during inference. We run inference both with and without prompts. For running inference without prompt, we generate 100,000 synthetic instructions using each trained model and check for the presence of secrets in the generations. The hyperparameters used for sampling are the same as those used in Section 4. For running inference with prompts, we utilize the text preceding the secret as the prompt and examine if the model’s completion, through greedy decoding, contains the secret. Table 3 presents the leakage of injected secrets. For models trained without DP, the loss ranks of all inserted canaries are alarmingly low, indicating that these models memorize the injected secrets. Furthermore, non-private models leak the secrets when canaries are repeated for 10 or 100 times, whereas private models demonstrate no such leakage.